A) Apply Kirchoff's Junction rule to the point E i

A) Apply Kirchoff's Junction rule to the point E in the given circuit

B) Apply Kirchoff's loop rule to loop EBCDE.

C) Apply Kirchoff's loop rule to loop FABEF
Answer #1

A)

from junction rule


I2 + I3 = I1


B)


from kirchoff loop rule


for the loop EBCDE

+9*I2 - 18 + 1*I2 - 15*I3 + 5 - 1*I3 - 14*I3 = 0

10I2 - 13 - 30I3 = 0


C)


from kirchoff loop rule


for the loop FABEF


-10 - 1*I1 - I1*17 - 1*I2 + 18 - 9*I2 - 2*I1 = 0

8 - 20*I1 - 10*I2 = 0


solving the 3 equations


I1 = 0.172 A

I2 = 0.454 A


I3 = -0.282 A